Use traditional method. Test statistic = 2.321, n1= 35 n2 = 35, α = 0.05, right tail.state whether the null hypothesis should be rejected or not. (Reject H0 / Do not Reject H0). 


1
Expert's answer
2021-01-21T19:20:52-0500

Degrees of freedom: df=n1+n2-2=35+35-2=68.

P-value: using EXCEL function =T.DIST.RT(2.321,68), we have p=0.0116.

Since the P-value is less than 0.05, reject the null hypothesis.